OpenStudy (anonymous):

If M (6,8) is the midpoint of line segment AB, and if A has coordinates (2,3) , find the coordinates of B.

OpenStudy (anonymous):

If Coordinates of B are (x,y), you know the mid point is calculated as: (2+x)/2=6 and (3+y)/2=8

OpenStudy (anonymous):

@nfarrar1313 can you find x and y with these equations?

OpenStudy (anonymous):

ok this one is easy.. just use midpoint theorem let A=(2,3)=(x,y) B=(c,d) let co-ordinates of midpoint be (e,f) by midpoint theorem you have e=(x+c)/2 so 6=(2+c)/2 hence c=10 similarly f=(y+d)/2 so 8=(3+d)/2 hence d=13 so here is your answer (10,13)